if mod of x is less than one and f(x) is equal to 1+x + x^2 and so on up to infinity then f (x) equal to

Grade:12

1 Answers

Aditya Gupta
5 years ago
Dear student, by the formula for infinite GP, we have
1+x+x^2+.....= 1/(1-x)= (1-x)^(-1)
If you have further doubts you can ask